# Weak Existence and Uniqueness for Super-Brownian Motion with Irregular Drift

### v1

@@ 4 Proof Outline
In this section, we outline the proof of Theorem 1.1 and, in particular, motivate the duality relation, which is the main tool in our analysis. Before we outline our arguments in more detail, we define what we understand under a weak solution to (1.1).
@@
Both, the proof of weak uniqueness and existence in Theorem 1.1 are based on a duality approach.
@@
$h_{\infty}\equiv\langle\nu_{1},1\rangle\text{ and }\nu_{2}\equiv 0,$ (4.2)
@@
$\exp(-\langle\phi,X_{t}\rangle)-\int_{0}^{t}\exp(-\langle\phi,X_{s}\rangle)\Big{(}-\langle\frac{1}{2}\Delta\phi,X_{s}\rangle+\frac{1}{2}\langle\phi^{2},X_{s}\rangle-\langle\phi,h(X_{s})\rangle\Big{)}ds$ (4.3)
@@
$\exp(-\langle Y_{t},\psi\rangle)-\int_{0}^{t}\exp(-\langle Y_{s},\psi\rangle)\Big{(}-\langle\frac{1}{2}\Delta Y_{s},\psi\rangle+\frac{1}{2}\langle Y_{s}^{2},\psi\rangle-\langle Y_{s},h(\psi)\rangle\Big{)}ds,$ (4.4)
@@
by [11, Theorem 4.4.11]. This would allow us to infer that the one-dimensional laws are unique and thus by standard arguments the finite dimensional laws of $X$ are unique as well. Consider the SPDE
@@
$h(x)=\int_{0}^{\infty}(1-\operatorname{e}^{-\lambda x})\nu^{1}(d\lambda)+\int_{0}^{\infty}(1+\operatorname{e}^{-\lambda x})\nu^{2}(d\lambda)$ (4.10)
@@
$\nu_{n}^{1}=\nu^{1}_{|(0,n]}+d_{1}\delta_{n},\quad\nu_{n}^{2}=\nu^{2}_{|(0,n]}+d_{2}\delta_{n}$
@@
Since $d_{1},d_{2}\geq 0$ this does in particular not cover the regime $b_{0}\geq 0$ and $b_{1}\leq 0$. Introducing an additional parameter $a\in\mathbb{R}$ and setting
$h_{\infty}(x)=2d_{2}\mathbbm{1}_{x=0}+(d_{1}+d_{2})\mathbbm{1}_{x>0}+\langle\nu^{1}+\nu^{2},1\rangle+a$ (4.12)
@@
It is easily verified, that in this case $d_{1},d_{2}\geq 0$ and by plugging these definitions into (4.12) we recover (1.4).
@@
There is not necessarily a unique choice for $\nu^{1},\nu^{2}$ and $d_{1}$, $d_{2}$ and $a$ that represent a given function $h$.
In order to handle this class of drift function we need to modify the duality function again. In this case we will see in Section 6 that the approximate duality relation is given by
@@
where $Y^{n}$ and $J^{n}$ are defined as in (4.11) with $d_{1},d_{2}$ as in (A). The relation (4.13) can again be easily motivated by a formal application of Ito’s formula. Hereby, we want to point out that $a$ may be negative. This means, one of our main tasks is to show that the expectation on right hand side of (4.13) is well defined. As before, we are able to construct the limits of $Y^{n}$ and $J^{n}$ and thus do away with the limit in (4.13). These considerations lead us to the following proposition, which we prove in Section 6
@@
Assume that $(X_{t})_{t\geq 0}$ is a solution to (1.1) with $X_{0}\sim\eta\in\mathcal{P}_{p}(\mathcal{C}_{rap}^{+})$, $p\geq 2$ and such that (1.5) holds. Let $Y_{0}=\mu$ for $\mu\in\mathcal{M}_{F}^{+}$. Then, there exists $T_{0}>0$ such that there exist processes $(J_{t})_{t\geq 0}$ and $(Y_{t})_{t\geq 0}$, taking values in $\mathbb{N}$ and $\mathcal{M}_{reg}^{+}$, respectively, that are independent of $X$ and such that
@@
Note that $(Y,J)$ are the limits of $(J^{n},Y^{n})$ in (4.13) that we discussed above with $d_{1},d_{2}$ as in (A).
Finally, when $\nu_{2}\equiv 0$ and $h_{\infty}\equiv\langle\nu_{1},1\rangle$ we can use Proposition 4.3 to show that the cozero set of solutions to (1.1) has finite Lebesgue measure. That is, we obtain the following result.
Theorem 4.5.
Suppose that almost surely, $X_{0}=f\in\mathcal{C}_{c}^{+}(\mathbb{R})$. And let $(X_{t})_{t\geq 0}$ be the solution to (1.1) with $\nu_{n}^{2}\equiv 0$ and $b_{0}=0$. Then it follows that there exists $T_{0}>0$ such that for all $t\in[0,T_{0}]$ we have
$\mathbb{P}(\mathrm{Leb}(\{x\in\mathbb{R}\colon X_{t}(x)>0\})<\infty)=1.$
@@ 5.0.2 Completely Monotone Functions
$h(x)=h_{1}(x)=\int_{0}^{\infty}\operatorname{e}^{-\lambda x}(\nu_{2}-\nu_{1})(d\lambda)$ (5.1)
where $\langle\nu_{2}-\nu_{1},1\rangle\geq 0$. By Bernsteins theorem on completely monotone functions (see [4]) all functions in the span of completely monotone functions that are non-negative at zero take the form (5.1). Recall, that a completely monotone function $f$ on $0,\infty)$ is continuous everywhere and in $\mathcal{C}^{\infty}((0,\infty))$ and satisfies
@@ 5.0.3 Discontinuous Drift
However, this can be proven along the same lines as the proof of [3, Lemma 1.3]. That is,by [24, Theorem VI.1.7] we get that
@@
where $L_{t}^{0}(y)$ is the local time of $y$ at 0. By the same arguments as in the proof of [24, Proposition XI.1.5] we get that $L_{t}^{0}(y)\equiv 0$. This finishes the proof of $(i)$.
@@
Finally, we prove $(iii)$. Assume that there exists a weak solution $x$ to (5.5) and note that this solution has to be non-negative by a slight modification of [16, Theorem 1.4 ]. Then, similar to above we get that
@@
The proof is similar to the proof of [3, Lemma 1.2]. Fix $c_{1},c_{2}\in(0,1)$ with $c_{1}\neq c_{2}$ and denote by $X^{1},X^{2}$ the solution to (5.7) with drift $h_{c_{1},1}$ and $h_{c_{2},1}$, respectively. We prove the claim by contradiction. To this end assume that $X^{1}\overset{d}{=}X^{2}$. Then it follows along the same lines as in the prove of [3, Lemma 1.2] that
@@
by a comparison argument, which follows along similar lines to the proof of [3, Lemma 1.2]. Let $X$ be the unique weak solution to
@@
And that $X^{i}$ as well as $X^{1}$ can be constructed as weak limits of an SPDE with Lipschitz continuous coefficients. Thus, by [25, Corollary 2.4] and weak uniqueness we get that $X$ stochastically dominates $X^{i}$, which implies
@@
However, using the duality formula for the Super Brownian motion with immigration given by (1.7) we get for $t>0$ that
@@ 6 Duality
$\nu([b,\infty])=(\nu^{1}+\nu^{2})([b,\infty))+d_{1}+d_{2},\quad b\in\overline{\mathbb{R}}_{+}.$
@@
$\mathbb{P}(M_{k}=i)=\frac{\nu^{i}((0,\infty))+d_{i}}{\nu((0,\infty])},\quad i\in\{1,2\},$ (6.1)
@@ 6.1 An Artificial Branching Process
The aim of this section is to derive bounds for $Y$, which only depend on $Y_{0}$ and $\nu((0,\infty])$. These bounds are crucial in order to prove that $Y$ only has finitely many jumps in each finite time interval and in order to show that all the terms in (4.14) have finite moments. To this end, we introduce an artificial branching mechanism that will allow us to keep track of the number of jumps of $Y$. This approach is inspired by the ideas from [3].
@@
$I_{t}(Y_{0})\subset\mathcal{U}=\cup_{k=1}^{\infty}\mathbbm{N}^{k}$
be the collections of labels of particles that are alive at time $t$. And denote by
@@
    $\mathfrak{N}^{(1)}(G_{1},G_{2},G_{3},G_{4})=\nu(G_{1})\int_{G_{2}}\int_{G_{3}}\int_{G_{4}}\mathbbm{1}_{Y_{t}^{(1)}(x)\leq r}dxdrdt,$
@@
  2. 6.2.
    To every particle $\mathcal{Z}^{(\alpha)}$ that was born at time $\tau_{\alpha}$ we associate a tuple $(Y^{\alpha},\mathfrak{M}^{\alpha})$, where $Y^{(\alpha)}_{t}=\mathcal{W}_{t-\tau_{\alpha}-}(0,0)\mathbbm{1}_{t>\tau_{\alpha}}$. Furthermore, $\mathfrak{M}^{(\alpha)}$ is a Poisson point process on $(0,\infty]\times(0,\infty)^{2}\times\mathbb{R}$ with intensity given by (6.6). Hereby, the processes $\mathfrak{M}^{(\alpha)}$ are independent. Moreover, we define the random measure $\mathfrak{N}^{(\alpha)}$ given by
    $\mathfrak{N}^{(\alpha)}(G_{1},G_{2},G_{3},G_{4})=\nu(G_{1})\int_{G_{2}}\int_{G_{3}}\int_{G_{4}}\mathbbm{1}_{Y_{t}^{(\alpha)}(x)\leq r}dxdrdt,$
@@
  3. 6.3.
    A particle $\mathcal{Z}^{(\alpha)}$, $\alpha\in\mathcal{U}$ gives birth to a new particles at time $\tau$, which is an arrival time of $\mathfrak{N}^{(\alpha)}$. If $\tau$ is the time of the $i$-th branching of $\mathcal{Z}^{(\alpha)}$ then the new particle will be labeled $\beta=(\alpha,i)$ and the parent keeps the label $\alpha$.
Recall that we denote by $I_{t}(Y_{0})$ the set of particles that are alive at time $t$. That is we have
$I_{t}(Y_{0})=\{\alpha\in\mathcal{U}\colon\langle Y_{t}^{(\alpha)},1\rangle>0\}.$
@@
First we prove the statement for $T_{0}>0$ sufficiently small. First, we define
$I_{t}^{0}=\{(1)\},\quad I_{t}^{i}=\{\alpha\in\mathcal{U}||\alpha|=i+1,\langle Y_{t}^{(\alpha)},1\rangle>0\},\quad i\geq 1$
@@
First, we prove (6.7) for the case $i=1$. The expected number of children of the particle $Z^{(1)}$ until time $t\in[0,T_{0}]$ is given by the expectation of $\mathfrak{N}^{(1)}_{t}=\mathfrak{N}^{(1)}((0,\infty],(0,t],(0,\infty),\mathbb{R})$, that is
@@
Since $(t,x)\mapsto V_{t}(Y_{0})(x)$ is integrable on $0,T]\times\mathbb{R}$ for all $T>0$, we can find $T_{0}>0$ sufficiently small so that
@@
$\mathbb{E}[|I_{t}^{{i+1}}|]=\mathbb{E}\Big{[}\sum_{\alpha\in I_{t}^{i}}\mathfrak{N}^{(\alpha)}_{t}\Big{]}\leq\mathbb{E}\Big{[}\sum_{\alpha\in I_{t}^{i}}\tilde{\mathfrak{N}}^{(\alpha)}_{t}\Big{]}\leq\mathbb{E}[|I_{t}^{i}|]\nu((0,\infty])\int_{0}^{t}\int_{\mathbb{R}}\mathcal{W}_{s}(0,0)(x)dxds\leq 2^{-i}2^{-1},$
@@
Hereby, we used that the intensity of $\tilde{\mathfrak{N}}^{(\alpha)}$ is larger than the intensity of $\mathfrak{N}^{(\alpha)}((0,\infty],\cdot,(0,\infty),\mathbb{R})$ for each $\alpha$. Thus, we arrive at
$\mathbb{E}[|I_{t}|]\leq\sum_{i=0}^{\infty}\mathbb{E}[|I_{t}^{i}|]\leq 1+\sum_{i=1}^{\infty}2^{-i},$
@@
where $I_{t}(\mathcal{W})$ denotes the particles that are alive at time $t$ in the branching process described in 6.1.-6.3. where $Y^{(1)}$ is replaced by $\mathcal{W}(0,0)$. This simply follows from the fact that
@@
$\begin{split}&\mathbb{E}[|I_{T}(Y_{0})|]\leq\mathbb{E}[|I_{MT_{0}}(Y_{0})|]\\ &\leq\mathbb{E}[|I_{T_{0}}(Y_{0})|-1]\mathbb{E}[|I_{(M-1)T_{0}}(\mathcal{W})|]+\mathbb{E}[|I_{(M-1)T_{0}}(V_{T_{0}}(Y_{0}))|]\leq\ldots\leq C(T)\end{split}$
@@
Let $Y_{0}=\mu+m\delta_{x_{0}}$ with $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$. Then for every $\gamma>0$, there exists $T_{0}>0$ such that
$\mathbb{E}[\operatorname{e}^{\gamma|I_{t}(Y_{0})||}]<\infty,$ (6.8)
for all $t\leq T_{0}$ . Hereby, $T_{0}>0$ only depends on $\nu((0,\infty])$, $\langle\mu,1\rangle$ and $\gamma$.
@@
Fix $t\leq T_{0}$, where $T_{0}$ is sufficiently small, the precise value will be chosen later. Define $\eta_{t}^{i,k}$ for $i,k\in\mathbb{N}$ to be i.i.d. Poisson distributed random variables with parameter
$\lambda=\nu((0,\infty])\Big{(}\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds+\int_{0}^{t}\langle V_{s}(Y_{0}),1\rangle ds\Big{)}.$
@@
Then it easily follows that $|I_{t}^{i}|$ is stochastically dominated by $Z_{i}^{t}$ for each $i\geq 1$. Thus, we can also infer that
$|I_{t}|=\sum_{i=0}^{\infty}|I_{t}^{i}|$
@@
The random variable $Z^{t}$ is nothing else, but the total progeny of $(Z^{t}_{n})_{n\geq 0}$ and its law can be made explicit by using the Otter-Dwass formula (see [9]), whenever $\lambda\leq 1$. Thus $T_{0}$ must be small enough such that $\lambda\leq 1$. In our setting, we get
@@
which is the Borel-Tanner distribution. Using Stirling’s approximation, we thus get that
@@
whenever
$\gamma\leq\lambda-1-\log(\lambda).$ (6.9)
When $\lambda\rightarrow 0$, we get that $\lambda-1-\log(\lambda)\rightarrow\infty$.
If we choose $T_{0}>0$ small enough, we thus get that (6.9) is satisfied and we obtain that
@@
Note that the function $x\mapsto\operatorname{e}^{\gamma x}$ is non-decreasing. Thus, the fact that $|I_{t}|$ is stochastically dominated by $Z^{t}$ implies that
$\mathbb{E}[\operatorname{e}^{\gamma|I_{t}|}]\leq\mathbb{E}[\operatorname{e}^{\gamma Z^{t}}]<\infty.$
This completes the proof. ∎
Now, we need to relate the branching process to the process $Y$. Since the birth times of new particles happen with higher frequency than the jumps of $t\mapsto\langle Y_{t},1\rangle$, we aim to control its time integral, which is sufficient for our arguments. Recall that $(S_{i})_{i=1}^{\infty}$ is the sequence of independent $\exp(1)$-distributed random variables from the construction of $Y$ as defined in (6.2). We now define random variables $(\bar{R}_{i})_{i=1}^{\infty}$ as follows
@@
$\langle V_{s}(Y_{0}),1\rangle\leq\langle V_{t}(Y_{0}),1\rangle\quad 0<t\leq s,$ (6.15)
we therefore arrive at
$\begin{split}&\langle Y_{\bar{T}_{i}+s},1\rangle\leq\langle V_{s+\bar{R}_{i}}(Y_{0}),1\rangle+\sum_{k=1}^{i}\langle\mathcal{W}_{s+\bar{R}_{i}-\bar{R}_{k}}(0,0),1\rangle\\ &\leq\langle\hat{Y}_{\bar{R}_{i}+s},1\rangle\quad a.s.,\end{split}$
for all $0<s<\bar{T}_{i+1}$. Hereby, (6.14) follows from Proposition 3.7 and (6.15) follows from from integrating (3.5).
Now, we show
@@
Using (6.12) for $i+1$ and using that $\bar{R}_{i}\leq\bar{T}_{i}$ we get

### v2

@@ 4 Proof Outline
In this section, we outline the proof of Theorem 1.1 and, in particular, motivate the duality relation, which is the main tool in our analysis. Before we outline our arguments in more detail, we give a definition of weak solution to (1.1).
@@
The proofs of weak uniqueness and of particular cases of weak existence in Theorem 1.1 are based on a duality approach.
@@
$h_{\infty}\equiv\langle\nu^{1},1\rangle\text{ and }\nu^{2}\equiv 0,$ (4.2)
@@
$\exp(-\langle\phi,X_{t}\rangle)-\int_{0}^{t}\exp(-\langle\phi,X_{s}\rangle)\Big(-\langle\frac{1}{2}\Delta\phi,X_{s}\rangle+\frac{1}{2}\langle\phi^{2},X_{s}\rangle-\langle\phi,h(X_{s})\rangle\Big)ds,\quad t\geq 0,$ (4.3)
@@
$\exp(-\langle Y_{t},\psi\rangle)-\int_{0}^{t}\exp(-\langle Y_{s},\psi\rangle)\Big(-\langle\frac{1}{2}\Delta Y_{s},\psi\rangle+\frac{1}{2}\langle Y_{s}^{2},\psi\rangle-\langle Y_{s},h(\psi)\rangle\Big)ds,\quad t\geq 0,$ (4.4)
@@
by [15, Theorem 4.4.11]. This would allow us to infer that the one-dimensional laws are unique and thus by standard arguments the finite dimensional laws of $X$ are unique as well. Consider the SPDE
@@
$h(x)=\int_{0}^{\infty}(1-\operatorname{e}^{-\lambda x})\nu^{1}(d\lambda)+\int_{0}^{\infty}(1+\operatorname{e}^{-\lambda x})\nu^{2}(d\lambda),$ (4.10)
@@
$\nu_{n}^{1}=\nu^{1}_{|(0,n]}+d_{1}\delta_{n},\quad\nu_{n}^{2}=\nu^{2}_{|(0,n]}+d_{2}\delta_{n},$
@@
Since $d_{1},d_{2}\geq 0$, for the case of $\nu^{1},\nu^{2}\equiv 0$ this, for example, does not cover the regime of $b_{0}>0$ and $b_{1}\leq 0$. Introducing an additional parameter $a\in\mathbb{R}$ and setting
$h_{\infty}(x)=2d_{2}\mathbbm{1}_{x=0}+(d_{1}+d_{2})\mathbbm{1}_{x>0}+\langle\nu^{1}+\nu^{2},1\rangle+a,$ (4.12)
@@
It is easily verified, that in this case of $d_{1},d_{2}\geq 0$, by plugging these definitions into (4.12) we recover (1.4).
@@
There is not necessarily a unique choice of $\nu^{1},\nu^{2}$ and $d_{1}$, $d_{2}$ and $a$ that allows to represent a given function $h$.
In order to handle this class of drift functions we need to modify the duality function again. In this case we will see in Section 6 that the approximate duality relation is given by
@@
where $Y^{n}$ and $J^{n}$ are defined as in (4.11) with $d_{1},d_{2}$ as in (A). The relation (4.13) can again be easily motivated by a formal application of Ito’s formula. Hereby, we want to point out that $a$ may be negative. This means, one of our main tasks is to show that the expectation on right hand side of (4.13) is well defined. As before, we are able to construct the limits of $Y^{n}$ and $J^{n}$ and thus do away with the limit in (4.13). These considerations lead us to the following proposition, which we will prove in Section 6
@@
Assume that $(X_{t})_{t\geq 0}$ is a solution to (1.1) with $X_{0}\sim\eta\in\mathcal{P}_{p}(\mathcal{C}_{rap}^{+})$, $p\geq 2$ and such that (1.5) holds. Let $Y_{0}=\mu$ for $\mu\in\mathcal{M}_{F}^{+}$. Then, there exists $T_{0}>0$ such that there exist processes $(J_{t})_{t\geq 0}$ and $(Y_{t})_{t\geq 0}$, taking values in $\mathbb{N}\cup\{0\}$ and $\mathcal{M}_{reg}^{+}$, respectively, that are independent of $X$ and such that
@@
Note that $(Y,J)$ are the limits of $(J^{n},Y^{n})$ in (4.13) that we discussed above with $d_{1},d_{2}$ as in (A). Finally, when $\nu^{2}\equiv 0$ and $h_{\infty}\equiv\langle\nu^{1},1\rangle$ we can use Proposition 4.3 to prove Theorem 1.2.
@@ 5.0.2 Completely Monotone Functions
$h(x)=h_{1}(x)=\int_{0}^{\infty}\operatorname{e}^{-\lambda x}(\nu^{2}-\nu^{1})(d\lambda)$ (5.1)
where $\langle\nu^{2}-\nu^{1},1\rangle\geq 0$. By Bernsteins theorem on completely monotone functions (see [5]) all functions in the span of completely monotone functions that are non-negative at zero take the form (5.1). Recall, that a completely monotone function $f$ on $0,\infty)$ is continuous everywhere and in $\mathcal{C}^{\infty}((0,\infty))$ and satisfies
@@ 5.0.3 Discontinuous Drift
However, this can be proven along the same lines as the proof of [4, Lemma 1.3]. That is, by [33, Theorem VI.1.7] we get that
@@
where $L_{t}^{0}(y)$ is the local time of $y$ at 0. By the same arguments as in the proof of [33, Proposition XI.1.5] we get that $L_{t}^{0}(y)\equiv 0$. This finishes the proof of $(i)$.
@@
Finally, we prove $(iii)$. Assume that there exists a weak solution $x$ to (5.5) and note that this solution has to be non-negative by a slight modification of [23, Theorem 1.4]. Then, similar to above we get that
@@
The proof is similar to the proof of [4, Lemma 1.2]. Fix $c_{1},c_{2}\in(0,1)$ with $c_{1}\neq c_{2}$ and denote by $X^{1},X^{2}$ the solution to (5.7) with drift $h_{c_{1},1}$ and $h_{c_{2},1}$, respectively. We prove the claim by contradiction. To this end assume that $X^{1}\overset{d}{=}X^{2}$. Then it follows along the same lines as in the prove of [4, Lemma 1.2] that
@@
by a comparison argument, which follows along similar lines to the proof of [4, Lemma 1.2]. Let $X$ be the unique weak solution to
@@
And that $X^{i}$ as well as $X^{1}$ can be constructed as weak limits of an SPDE with Lipschitz continuous coefficients. Thus, by [35, Corollary 2.4] and weak uniqueness we get that $X$ stochastically dominates $X^{i}$, which implies
@@
However, using the duality formula for the Super-Brownian motion with immigration given by (1.12) we get for $t>0$ that
@@ 6 Duality
$\nu([b,\infty])=(\nu^{1}+\nu^{2})([b,\infty))+d_{1}+d_{2},\quad b\in\mathbb{R}_{+}.$ (6.1)
@@
$\mathbb{P}(M_{1}=i)=\frac{\nu^{i}((0,\infty))+d_{i}}{\nu((0,\infty])},\quad i\in\{1,2\},$ (6.2)
@@ 6.1 An Artificial Branching Process
The aim of this section is to derive bounds for $Y$, which only depend on $Y_{0}$ and $\nu((0,\infty])$. These bounds are crucial in order to prove that $Y$ only has finitely many jumps in each finite time interval and in order to show that all the terms in (4.14) have finite moments. To this end, we introduce an artificial branching mechanism that will allow us to control the number of jumps of $Y$. This approach is inspired by the ideas from [4].
@@
$I^{\nu}_{t}(Y_{0})\subset\mathcal{U}=\cup_{k=1}^{\infty}\mathbbm{N}^{k}$
be the collections of labels of particles that are alive at time $t$. When there is no ambiguity we may suppress the superscript $\nu$ and just write $I_{t}(Y_{0})$ . And denote by
@@
    $\mathfrak{N}^{(1)}(G_{1},G_{2},G_{3},G_{4})=\nu(G_{1})\int_{G_{2}}\int_{G_{3}}\int_{G_{4}}\mathbbm{1}_{r\leq Y_{t}^{(1)}(x)}dxdrdt,$
@@
  2. 6(ii)
    To every particle $\mathcal{Z}^{(\alpha)}$ that was born at time $\tau_{\alpha}$ we associate a tuple $(Y^{(\alpha)},\mathfrak{M}^{(\alpha)})$, where $Y^{(\alpha)}_{t}=\mathcal{W}_{t-\tau_{\alpha}-}(0,0)\mathbbm{1}_{t>\tau_{\alpha}}$. Furthermore, $\mathfrak{M}^{(\alpha)}$ is a Poisson point process on $(0,\infty]\times(0,\infty)^{2}\times\mathbb{R}$ with intensity given by (6.7). Hereby, the processes $\mathfrak{M}^{(\alpha)}$ are independent. Moreover, we define the random measure $\mathfrak{N}^{(\alpha)}$ given by
    $\mathfrak{N}^{(\alpha)}(G_{1},G_{2},G_{3},G_{4})=\nu(G_{1})\int_{G_{2}}\int_{G_{3}}\int_{G_{4}}\mathbbm{1}_{r\leq Y_{t}^{(\alpha)}(x)}dxdrdt,$
@@
  3. 6(iii)
    A particle $\mathcal{Z}^{(\alpha)}$, $\alpha\in\mathcal{U}$ gives births to new particles at arrival times of $\mathfrak{M}^{(\alpha)}$. For the $i$-th branching of $\mathcal{Z}^{(\alpha)}$, the new particle will be labeled $\beta=(\alpha,i)$ and the parent keeps the label $\alpha$.
Recall that we denote by $I_{t}(Y_{0})$ the set of particles that are alive at time $t$. That is, we have
$I_{t}(Y_{0})=\{\alpha\in\mathcal{U}\colon\langle Y_{t}^{(\alpha)},1\rangle>0\}.$ (6.8)
@@
Define
$I_{t}^{0}=\{(1)\},\quad I_{t}^{i}=\{\alpha\in\mathcal{U}||\alpha|=i+1,\langle Y_{t}^{(\alpha)},1\rangle>0\},\quad i\geq 1,$
@@
First, we prove (6.9) for the case $i=1$. The expected number of children of the particle $Z^{(1)}$ until time $t\in[0,T_{0}]$ is given by the expectation of $\mathfrak{N}^{(1)}_{t}=\mathfrak{N}^{(1)}((0,\infty]\times(0,t]\times(0,\infty)\times\mathbb{R})$, that is
@@
Since, by Lemma 3.2, $(t,x)\mapsto V_{t}(Y_{0})(x)$ is integrable on $0,T]\times\mathbb{R}$ for all $T>0$, we can find $T_{0}>0$ sufficiently small so that
@@
$\begin{split}\mathbb{E}[|I_{t}^{{i+1}}|]&=\mathbb{E}\Big[\sum_{\alpha\in I_{t}^{i}}\mathfrak{N}^{(\alpha)}((0,\infty]\times(0,t]\times(0,\infty)\times\mathbb{R})\Big]\\ &\leq\mathbb{E}\Big[\sum_{\alpha\in I_{t}^{i}}\tilde{\mathfrak{N}}^{(\alpha)}_{t}\Big]\leq\mathbb{E}[|I_{t}^{i}|]\nu((0,\infty])\int_{0}^{t}\int_{\mathbb{R}}\mathcal{W}_{s}(0,0)(x)dxds\leq 2^{-i}2^{-1},\end{split}$
@@
Hereby, we used that the intensity of $\tilde{\mathfrak{N}}^{(\alpha)}$ bounds the intensity of $\mathfrak{N}^{(\alpha)}((0,\infty],dt,(0,\infty),\mathbb{R})$ from above for each $\alpha$. Thus, we arrive at
$\mathbb{E}[|I_{t}(Y_{0})|]=\sum_{i=0}^{\infty}\mathbb{E}[|I_{t}^{i}|]\leq 1+\sum_{i=1}^{\infty}2^{-i},$
@@
where $I_{t}(\mathcal{W})$ denotes the particles that are alive at time $t$ in the branching process described in 6(i)-6(iii), where $Y^{(1)}$ is replaced by $\mathcal{W}(0,0)$. This simply follows from the fact that
@@
$\begin{split}&\mathbb{E}[|I_{T}(Y_{0})|]\leq\mathbb{E}[|I_{MT_{0}}(Y_{0})|]\\ &\leq\mathbb{E}[|I_{T_{0}}(Y_{0})|-1]\mathbb{E}[|I_{(M-1)T_{0}}(\mathcal{W})|]+\mathbb{E}[|I_{(M-1)T_{0}}(V_{T_{0}}(Y_{0}))|]\leq\ldots\leq C(T).\end{split}$
@@
Let $Y_{0}=\mu+m\delta_{x_{0}}$ with $\mu\in\mathcal{M}_{F}^{+}$, $x_{0}\in\mathbb{R}$ and $m\in\mathbb{N}\cup\{\infty\}$. Furthermore, let $I_{t}(Y_{0})$ be defined as in (6.8). Then for every $\gamma>0$, $K>0$, there exists $T_{0}>0$ such that
$\sup_{\nu:\nu((0,\infty])\leq K}\mathbb{E}[\operatorname{e}^{\gamma|I_{t}(Y_{0})||}]<\infty,$ (6.10)
for all $t\leq T_{0}$ . Hereby, $T_{0}>0$ only depends on $K$, $\langle\mu,1\rangle$ and $\gamma$.
@@
Fix arbitrary $\gamma>0$, $K>0$ and $t\leq T_{0}$, where $T_{0}$ is sufficiently small, the precise value will be chosen later. Define $\eta_{t}^{i,k}$ for $i,k\in\mathbb{N}$ to be i.i.d. Poisson distributed random variables with parameter
$\lambda=K\Big(\int_{0}^{t}\langle\mathcal{W}_{s}(0,0),1\rangle ds+\int_{0}^{t}\langle V_{s}(Y_{0}),1\rangle ds\Big).$ (6.11)
@@
Let $\nu$ be a measure such that $\nu((0,\infty])\leq K$. Then it easily follows that $|I_{t}^{i}|$ is stochastically dominated by $Z_{i}^{t}$ for each $i\geq 1$. Thus, we can also infer that
$|I_{t}(Y_{0})|=\sum_{i=0}^{\infty}|I_{t}^{i}|$
@@
The random variable $Z^{t}$ is nothing else, but the total progeny of $(Z^{t}_{n})_{n\geq 0}$ and its law can be made explicit by using the Otter-Dwass formula (see [13]), whenever $\lambda\leq 1$. By (6.11) we can choose $T_{0}$ small enough such that
$\lambda\leq 1,\quad\log(\lambda)<\gamma-1,\quad\forall t\leq T_{0}.$ (6.12)
Then we get
@@
that is $Z^{t}$ has the Borel-Tanner distribution. Using Stirling’s approximation, we thus get that
@@
by (6.12). Therefore, we obtain that
@@
Note that the function $x\mapsto\operatorname{e}^{\gamma x}$ is non-decreasing. Thus, the fact that $|I_{t}(Y_{0})|$ is stochastically dominated by $Z^{t}$ implies that
$\mathbb{E}[\operatorname{e}^{\gamma|I_{t}(Y_{0})|}]\leq\mathbb{E}[\operatorname{e}^{\gamma Z^{t}}]<\infty.$
Since $\nu$ such that $\nu((0,\infty])\leq K$ was arbitrary, this completes the proof. ∎
Now, we need to relate the branching process to the process $Y$. Since the birth times of new particles in the branching process happen with higher rate than the jumps of $t\mapsto\langle Y_{t},1\rangle$, we aim to control $\int_{0}^{t}\langle Y_{s},1\rangle ds$, which is sufficient for our arguments. Recall that $(S_{i})_{i=1}^{\infty}$ is the sequence of independent $\exp(1)$-distributed random variables used for the construction of $Y$ as defined in the beginning of Section 6. We now define random variables $(\bar{R}_{i})_{i=1}^{\infty}$ as follows
@@
$\langle V_{s}(Y_{0}),1\rangle\leq\langle V_{t}(Y_{0}),1\rangle\quad 0<t\leq s.$ (6.18)
Hereby, (6.17) follows from Proposition 3.7 and (6.18) follows from integrating (3.5). We arrive at
$\begin{split}&\langle Y_{\bar{T}_{i}+s},1\rangle\leq\langle V_{s+\bar{R}_{i}}(Y_{0}),1\rangle+\sum_{k=1}^{i}\langle\mathcal{W}_{s+\bar{R}_{i}-\bar{R}_{k}}(0,0),1\rangle\\ &\leq\langle\hat{Y}_{\bar{R}_{i}+s},1\rangle\quad a.s.,\end{split}$ (6.19)
for all $0<s<\bar{T}_{i+1}$.
Now, we will show
@@
Using (6.15) for $i+1$, which follows from (6.19), and using that $\bar{R}_{i}\leq\bar{T}_{i}$, we get